OpenStudy (anonymous):

Take the derivative of y=x^(2x)

hartnn (hartnn):

logarithmic differentiation

OpenStudy (anonymous):

Is the answer 2x(x)^(2x-1) *(2)

OpenStudy (anonymous):

ah, no then

hartnn (hartnn):

yep, since you have exponent as variable, you'll have to take log on both sides first

hartnn (hartnn):

y = x^(2x) ln y = ...?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I forgot all these rules :o I think the bases switch

OpenStudy (anonymous):

maybe 2x(x) ?

OpenStudy (anonymous):

I need to look over these logs >_<

hartnn (hartnn):

\(\Large \ln a^b = b \ln a\)

OpenStudy (anonymous):

2xlnx then?

hartnn (hartnn):

yesss ln y = 2x ln x now differentiate both sides w.r.t x

OpenStudy (anonymous):

ok

hartnn (hartnn):

chain rule on left side product rule on right side

OpenStudy (anonymous):

ok so 1/x on the left?

hartnn (hartnn):

there's no 'x' on left :P

OpenStudy (anonymous):

hmm wait

OpenStudy (anonymous):

oh I'm hopeless with logs

OpenStudy (anonymous):

I'm stuck

hartnn (hartnn):

\(\Large d/dx (\ln y) = (1/y) d/dx (y) = (1/y)(dy/dx)\)

OpenStudy (anonymous):

So this rule applies, it's just not x, it's y

hartnn (hartnn):

got that ? and what about right side ?

OpenStudy (anonymous):

doing it

OpenStudy (anonymous):

(2)(lnx)+(2x)(1/x)

OpenStudy (anonymous):

which simplifies to 2lnx+2 if it's correct??

hartnn (hartnn):

correct!

hartnn (hartnn):

now just isolate dy/dx

hartnn (hartnn):

by multiplying y on both sides

hartnn (hartnn):

and then plugging in y = x^(2x)

OpenStudy (anonymous):

just so it's clear: currently (1/y)(dy/dx)=2lnx+2 ?

hartnn (hartnn):

correct

OpenStudy (anonymous):

ok then

hartnn (hartnn):

to isolate dy/dx ,what u need to do ??

OpenStudy (anonymous):

multiply by y

OpenStudy (anonymous):

dy/dx=y(2lnx+2)

OpenStudy (anonymous):

y=x^2x

hartnn (hartnn):

yes., go on, and factor out the 2 too

OpenStudy (anonymous):

so dy/dx=(x^2x)(2lnx+2)=

hartnn (hartnn):

factor the 2, and you'll be done :)

OpenStudy (anonymous):

(x^2x)(2)(lnx+1)

hartnn (hartnn):

thats it! correct :)

OpenStudy (anonymous):

whew, thanks :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

are logs very common in upper level calculus?

hartnn (hartnn):

not very common, but you will see them here and there.....always good to know them properly.

OpenStudy (anonymous):

ok, thanks for the advice